OpenStudy (anonymous):

The total cost to park a car in a parking garage that has an entry fee of $4 can be determined by using the equation c = 2h + 4 where c is the total cost and h is the number of hours the car is in the garage. The garage informed Robert that the entry fee had changed and that his bill for parking his car for 6 hours is $20. Which equation models the cost to park his car for h hours with the new entry fee? A. c = 2h + 20 B. c = h + 6 C. c = 2h + 8 D. c = h + 14

OpenStudy (anonymous):

if only entry fee changed, then 2h will still remain constant so 20=2*6+entry fee find entry fee.

OpenStudy (triciaal):

c = 2h + fee h stays the same 6 when fee = 4 cost = 2(6)+ 4 = 16 when cost = 20 = 16 + (20-16) = 2h + 4 + 4 = 2h + 8 new cost, c = 2h + 8